A piece of cork is floating on water in a small tank. The cork oscillates up and down vertically when small ripples pass over the surface of water. The velocity of the ripples being 0.21 m/s , wave length 15 mm and amplitude 5 mm, the maximum velocity of the piece of cork is 

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
We know that in wave motion , maximum particle velocity is given by , vmax​=ωA=2πfA,
whereω=angular frequency ,
f=linear frequency ,
A=amplitude ,
given - wave velocity ,u=0.21m/s,λ=15mm=0.015m,A=5mm=0.005m,
by , f=u/λ=0.21/0.015Hz ,
therefore maximum particle (cork) velocity ,
vmax​=2×(22/7)×(0.21/0.015)×0.005=0.44m/s